Subtonic:

Use the inequality to answer Parts 1-3. −3(x−2)≤d Part 1: Solve the inequality. Leave answer in terms of a whole number or reduced improper fraction. Part 2: Write a verbal statement describing the solution to the inequality. Part 3: Verify your solution to the inequality using two elements of the solution set.

Subtonic:

so this is the wrong website

Zeldris:

Use a word processing program or handwrite your responses to Parts 1-3. Turn in all three responses.

Joe348:

Part A: Multiply both sides by -1 \[3(x-2)\ge-d\] Divide both sides by 3 \[x-2\ge-d/3\] Add 2 to both sides \[x \ge-d/3 +2\]

Subtonic:

thanks joe

Joe348:

By any chance, you have the solution set for part 3?

Subtonic:

nope. this is exactly how it was formatted when I got it

Joe348:

Okay well for part 3, all you have to do is find the number that's greater than that number and find the solution Can't figure out the rest but @jhonyy9 might help you with parts 2 and 3!😄

Joe348:

Well, do you have any options? Like multiple choice?

Subtonic:

no. this is why i got so confused i asked on 2 websites 💀

Subtonic:

ok how about this

Subtonic:

1 attachment